# A Semantics-Aware Hierarchical Self-Supervised Approach to Classification of Remote Sensing Images

### v1

@@ III-D Self-Supervised Hierarchical Consensus Learning
Let us return to the general case of $H$ hierarchical levels, considering projections from source levels $h$ to a target hierarchical level $\tilde{h}$. Each hierarchical level acts independently of the others. To ensure that the network is hierarchy-aware, an additional constraint based on the prediction associated to the target hierarchical level and the projected predictions from the other levels is defined. In particular, we take inspiration from the disagreement-based active learning field [25] to define the disagreement between the predictions calculated at each hierarchical level. For each level $\tilde{h}$, we consider a committee composed of all $H$ levels. All predictions and their projected versions can be viewed as “soft” votes, which define each committee member’s confidence. Then, we introduce a self-supervised loss that penalizes the committee members’ disagreement with the overall committee consensus, thus guiding the learning of data-driven relationships in $\mathbf{J}^{h,\tilde{h}}$ and providing further supervision to the network based on the predictions at the different levels.
@@ III-D1 Hierarchy-based Projected Predictions
To further simplify, hierarchy projection matrices can be defined directly in the log domain, i.e., we learn log joint probabilities $\mathbf{L}^{h,\tilde{h}}=\log{\hat{\mathbf{J}}^{h,\tilde{h}}}\in\mathbb{R}^{|\Omega^{h}|\times|\Omega^{\tilde{h}}|}$. Then, we compute $\log{\hat{\mathbf{M}}^{h,\tilde{h}}}$ as follows:
@@ III-D2 Semantics-Aware Hierarchical Consensus
where $\text{LSE}{(\mathbf{g}^{\tilde{h}}_{HC})}=\log{Z^{\tilde{h}}}$. While not directly supervised, HC can serve as an in-place replacement for the original classification modules during inference, providing a fully semantics-aware prediction in the downstream task.
III-D3 Self-Supervised Consensus Loss
After calculating the consensus, both the prediction at the target level and the projected predictions are compared with the HC, defining an additional loss based on the divergence from the HC. To this purpose, we employ the JSD, which is defined as follows:
@@ III-E Overall Loss Function
$\mathcal{L}=\mathcal{L}_{CCE}+\lambda_{HC}\mathcal{L}_{HC},$ (18)
where $\lambda_{HC}$ is a metaparameters tuning the contribution of the HC loss. It is important to note that $\mathcal{L}_{HC}$ does not require supervision, as it is simply a loss representing the divergence from the HC. Though the HC prediction is used within the penalty loss during the training phase, we consider the HC prediction at each hierarchical level as the final predictions for inference. This allows the architecture to leverage the semantic relationships learnt during the self-supervised training phase and to perform hierarchically consistent predictions.
@@ IV-B ELU Dataset
Figure 3: A semicircular hierarchical label tree representation of the Emilia LU dataset at the fine-grained level of the hierarchy. The leaves corresponding to the fine-grained classes are shown in green, while the nodes for intermediate classes are depicted in blue, and those for coarse-grained classes are shown in red. The root node is represented in black.
@@
To further validate the performance of the proposed methodology, we additionally considered two multitemporal multispectral datasets differing in terms of complexity in the hierarchical structure. The first dataset [^1] Land Use of Emilia (2020). Accessed on May $13^{\text{th}}$ $2025$. consists of a publicly available LU map of $2020$ over the Emilia Romagna region, Italy. The maps were defined considering the standard CORINE Land Cover (CLC) aggregation at three levels, with the third one further divided, resulting in an additional fourth level with an increased semantic detail. For the purpose of this analysis, we evaluated the hierarchical approach on three levels of hierarchy, consisting in the first two levels of the CLC and the additional finest level of detail, here referred to as coarse-grained, intermediate and fine-grained, respectively. The dataset has been built considering orthophotos and infrared images, and shows a minimum mapping unit of $0.16$ hectares, with a minimum of $7\text{m}$ meters for linear structures. Although it has a good positional accuracy of $5\text{m}$, the data may exhibit errors. Additionally, it represents LU classes, where the same Land-Cover (LC) class could be associated with different semantic classes, such as the fine-grained classes “Commercial airports”, “Sport flight airports and heliports”, and “Military airports”. We considered the Emilia Land Use (ELU) dataset corresponding to the Military Grid Reference System (MGRS) Sentinel-2 (S2) tile 32TPQ ($10980\times 10980$ pixels, for a total area of $110\times 110$ $km^{2}$). In the analyzed tile, the three hierarchical levels show $5$, $18$, and $80$ classes at the coarse-grained, intermediate, and fine-grained levels, respectively. Please note that such an elevated number of fine-grained classes is not commonly approached in Earth Observation (EO) classification problems, where the majority of studies focus on the intermediate or coarse-grained classification task. Nonetheless, this level of precision allows a better monitoring of the class of interest and the extraction of information-rich content. Due to the high number of classes and, consequently, the large size of the hierarchical tree, we refer to this hierarchy as a complex hierarchy. A semicircular tree representing the hierarchical structure of the dataset is shown in Fig. 3.
IV-C ESA HRLC CCI Dataset
The second multitemporal multispectral dataset [^2] ESA High-Resolution Land Cover Climate Change Initiative Project CEDA Archive site. Accessed on May $13^{\text{th}}$ $2025$. has been derived under the High Resolution Land Cover (HRLC)- Climate Change Initiative (CCI) project by European Space Agency (ESA) [30]. The dataset includes LC classes organized according to the LCCS nomenclature, following the classification at the finest level of details. In this case, the hierarchical structure is less rigid compared to the Emilia dataset, as the classes are less detailed and can almost be considered independent. In this analysis, we use the HRLC1010 maps, i.e., the $10\text{m}$ resolution maps for the year $2019$. The maps have been derived from machine learning approaches considering a joint approach based on multispectral (MS) S2 and Synthetic Aperture Radar (SAR) Sentinel-1 (S1) images [31]. The HRLC1010 maps have been validated by intercomparison with other available maps and with photointerpreted samples. For our analysis, we considered the dataset corresponding to the MGRS S2 tile 22KGV ($10980\times 10980$ pixels, for a total area of $110\times 110$ $km^{2}$) in Brazil, mostly due to the heterogeneous presence of different classes. Since the hierarchical structure is simpler compared to the Emilia LU dataset, we defined three hierarchical levels ad-hoc, based on the similarity of the classes at the finest level of detail. This led to the definition of $3$, 77, and $12$ classes for the different levels, corresponding to coarse-grained, intermediate, and fine-grained classes, respectively. The coarse-grained classes have been divided in “Woody vegetated areas”, “Herbaceous vegetated areas”, and “Non-vegetated areas”. The fine-grained classes have been aggregated at the intermediate level considering the seasonality (e.g., “Tree cover evergreen broadleaf” and “Tree cover deciduous broadleaf” are aggregated at intermediate level, while “Tree cover evergreen needleleaf” is still separated and aggregated at the coarse-grained level instead). Since the number of classes only shows a slight increment in the different layers, we refer to this hierarchy as a simple hierarchy.
IV-D S2 Preprocessing and Dataset Preparation
IV-D1 VHR dataset
@@ IV-D2 Multitemporal MS datasets
The input data consists of S2 MS satellite images provided in the Copernicus EO program. The S2 acquisitions considered have been preprocessed at Level-2A (L2A) and re-projected at $10\text{m}$/pixel resolution through a nearest neighbor approach. In the experiments, all the bands except the $60\text{m}$ resolution band were used. Due to the high-complexity of both classification problems, we collected two time series (TS) of S2 images, fully exploiting the temporal signature and allowing the detection of seasonal targets. The S2 TS considered for the Emilia Dataset consists of acquisitions ranging from January $1^{\text{st}}$ $2020$ to December $31^{\text{st}}$ $2020$, while the S2 TS for the HRLC-CCI dataset consists of TS ranging from January $1^{\text{st}}$ $2019$ to December $31^{\text{st}}$ $2019$. Both the S2 TSs have been processed to achieve more regular TSs, resulting in $12$ median monthly composites. This was achieved by computing the pixel-wise median for each band for each month. Before applying the median operator, we masked the samples associated with saturate or defective pixels, clouds, and cloud shadows. This approach allowed us to have regular, gap-free data in input to the DL models.
@@
To extract a balanced training set, we considered an approach based on the heterogeneity of the input training patches. In particular, for both datasets, we divided the initial S2 monthly composite images into non-overlapping patches of $9\times 9$ pixels. After the initial subdivision, we collected the patches and calculated the mode, extracting the dominant class of each patch analyzed. Then, the patches have been randomly subdivided into training, validation, and test sets, by considering a minimum number of $N$ representative patches for each class in the analyzed classification scheme, with $N$ fixed to $1000$ for both the ELU and HRLC-CCI datasets. In the case of a minoritarian class had fewer patches than $N$, $50\%$ of its total were assigned to the training set, $25\%$ to the validation set, and $25\%$ to the test set. In this way, we collected a small training dataset representative of the entirety of the classification scheme, corresponding to $\sim 1.35\%$ ($1.631.616$ pixels) and to $\sim 0.22\%$ ($268.800$ pixels) of the total number of patches in the tile for the Emilia Dataset and the HRLC-CCI dataset, respectively. The validation set size has been determined by considering $25\%$ of the number of samples allocated for each class for the training set, while the remainder of the samples (i.e., the entirety of the patches in the tile not belonging to the training and validation sets) have been used to validate the proposed approach in the test set.
IV-E Design of Experiments
Instead of validating the proposed hierarchical approach on a single backbone, we demonstrated its broad applicability by employing diverse backbone architectures representing different paradigms in a robust manner. All tests were conducted five times, with different random initializations of the weights of the architecture. For reproducibility, predefined seed values were used for the random number generator (RNG). The reported results include the values of Overall Accuracy (OA) for both tasks, while mean F1 Score (mF1) was also employed in the SITS MS classification task, along with the standard deviation (calculated as an unbiased estimator). For each architecture, we considered the following approaches: (i) the baseline with only the fine-grained classification through standard CCE minimization, (ii) the HAF method [18], (iii) the HRN module [20] and (iv) the proposed Semantics-Aware Hierarchical Consensus (SAHC). We removed the Granularity-Specific Block (GSB) block (two parameter-heavy convolutions and fully connected layers) from the HRN approach, maintaining only the hierarchical module. This is to ensure a fair comparison of architectures with the same number of parameters and focus solely on the hierarchical representation of the problem.
The different strategies are evaluated on the VHR scene classification task and the multitemporal MS semantic segmentation tasks. For the VHR task, we considered three different backbone architectures focusing on three different paradigms: (i) EfficientformerV2 S2 [33] for transformer-based architectures, (ii) MobileNetV2 [34] for CNN-based architectures, and (iii) LWGANet [35] for hybrid architectures. For the MS SITS classification task, we considered three different backbone architectures designed to process temporal information, also including hybrid architectures to further examine the effectiveness of the methodologies in utilizing contextual information: (i) the encoder part of a Transformer [36], (ii) an adaptation to the RS scenario of a Swin Transformer originally intended for super-resolution of videos [37], and (iii) a Convolutional Long Short-Term Memory (ConvLSTM) [38]. For a detailed description of the layer topology, please refer to the cited papers.
@@ IV-E Design of Experiments
For each backbone, we replaced the classification head with three hierarchical classification heads, each characterized by a different level of granularity. We adopted the AdamW optimizer for all the architectures and configurations. The metaparameters of each different configuration have been validated considering a Tree-structured Parzen Estimator (TPE) algorithm to efficiently explore the search space. The learning rate and the weight decay were sampled from log-uniform distributions $U_{\log}([10^{-6},10^{-3}])$. Considering the proposed SAHC approach, the weight $\lambda_{HC}$ associated to the HC loss has been sampled from a log-uniform distribution $U_{\log}([10^{-2},10^{2}])$, while the weights $\lambda^{h}$ have been set empirically to $0.3$, $0.2$, and $0.5$, to prioritize the fine-grained classification and reduce the influence of the intermediate and coarse-grained levels. To adjust the learning rate during training, we considered a cosine decay schedule. The weights associated to each model in the best performing epoch in terms of validation OA% and mF1% were used to conduct the inference on the test set, whose performance metrics are reported in the following section.
@@ V Experimental Results
The experiments focus on validating the proposed SAHC approach over different backbones tasks, and datasets, comparing it to different standard and hierarchical approaches. We first focus on the results on the specific tasks, i.e., scene classification on the VHR dataset and semantic segmentation of MS SITS. Then, we demonstrate the effectiveness of each component of the SAHC approach in an ablation study considering the ELU dataset. Finally, we analyze the trained hierarchy projection matrices to identify the key hierarchical relationships learnt by the proposed approach.

### v2

@@ III-D Self-Consistent Hierarchical Consensus Learning
Let us return to the general case of $H$ hierarchical levels, considering projections from source levels $h$ to a target hierarchical level $\tilde{h}$. Each hierarchical level acts independently of the others. To ensure that the network is hierarchy-aware, an additional constraint based on the prediction associated to the target hierarchical level and the projected predictions from the other levels is defined. In particular, we take inspiration from the disagreement-based active learning field [26] to define the disagreement between the predictions calculated at each hierarchical level. For each level $\tilde{h}$, we consider a committee composed of all $H$ levels. All predictions and their projected versions can be viewed as “soft” votes, which define each committee member’s confidence. Then, we introduce a self-consistent loss that penalizes the committee members’ disagreement with the overall committee consensus, thus guiding the learning of data-driven relationships in $\mathbf{J}^{h,\tilde{h}}$ and providing further supervision to the network based on the predictions at the different levels.
@@ III-D1 Hierarchy-based Projected Predictions
Hierarchy projection matrices can be defined directly in the log domain, i.e., we learn log joint probabilities $\mathbf{L}^{h,\tilde{h}}=\log{\hat{\mathbf{J}}^{h,\tilde{h}}}\in\mathbb{R}^{|\Omega^{h}|\times|\Omega^{\tilde{h}}|}$. Then, we compute $\log{\hat{\mathbf{M}}^{h,\tilde{h}}}$ as follows:
@@ III-D2 Semantics-Aware Hierarchical Consensus
where $\text{LSE}{(\mathbf{g}^{\tilde{h}}_{HC})}=\log{Z^{\tilde{h}}}$. This HC represents the consensus found by the voting committee and can serve as an in-place replacement for the original classification modules during inference, providing a fully semantics-aware prediction in the downstream task.
III-D3 Self-Consistent Consensus Loss
After calculating the consensus, both the prediction at the target level and the projected predictions are compared with the HC, defining an additional loss based on the divergence from the HC. To this purpose, we emply a JSD loss rather than a standard Kullback–Leibler divergence (KLD). While the KLD is widely used to measure the difference between two probability distributions, its asymmetry makes it better suited for scenarios where one distribution is a fixed “ground truth” or prior. In this case, all hierarchical levels provide dynamically trained “soft” votes, with no reference distribution. The symmetry in the JSD ensures balanced regularization between the projected predictions and the consensus, also providing stabler gradients due to its bounded nature. In the proposed framework, the JSD loss is computed as follows:
@@ III-D3 Self-Consistent Consensus Loss
Finally, we also apply full supervision on the consensus predictions at the different levels considering a standard CCE loss:
$\mathcal{L}_{CCE}^{HC}=\sum_{h=1}^{H}-\frac{1}{N}\sum_{n=1}^{N}\sum_{i=1}^{|\Omega^{h}|}\mathbbm{1}\left[y^{h}_{n}=\omega^{h}_{i}\right]\log\hat{P}^{h}_{HC}\left(\omega^{h}_{i}|\mathbf{x}_{n}\right).$ (18)
@@ III-E Overall Loss Function
$\mathcal{L}=\mathcal{L}_{CCE}+\mathcal{L}_{CCE}^{HC}+\lambda_{HC}\mathcal{L}_{HC},$ (19)
where $\lambda_{HC}$ is a metaparameters tuning the contribution of the HC loss. It is important to note that $\mathcal{L}_{HC}$ does not require direct supervision, as it is simply a loss representing the divergence from the HC. The HC prediction is used both within the penalty loss, and the additional supervised CCE loss during the training phase. We further consider the HC prediction at each hierarchical level as the final predictions for inference. This allows the architecture to leverage the semantic relationships learnt during the self-consistent training phase and to perform hierarchically consistent predictions.
@@ IV-B ELU Dataset
Figure 4: A semicircular hierarchical label tree representation of the ELU dataset at the fine-grained level of the hierarchy. The leaves corresponding to the fine-grained classes are shown in orange, while the nodes for intermediate classes are depicted in green, those for coarse-grained classes are shown in blue, and the nodes for coarse in brown. The root node is represented in black.
@@
To further validate the performance of the proposed methodology and demonstrate its generability on different land cover mapping tasks, we additionally considered a multitemporal multispectral dataset differing in terms of complexity in the hierarchical structure. This dataset [^1] Land Use of Emilia (2020). Accessed on April $9^{\text{th}}$ $2026$. consists of a publicly available Land-Use (LU) map of $2020$ over the Emilia Romagna region, Italy. The maps were defined considering the standard CLC aggregation at four levels, showcasing a highly detailed semantic dataset. For the purpose of this analysis, we evaluated the hierarchical approach on four levels of hierarchy, consisting in the first three levels of the CLC and the additional finest level of detail defined by the LULC, here referred to as macro, coarse-grained, intermediate and fine-grained, respectively. The dataset has been built considering orthophotos and infrared images, and shows a minimum mapping unit of $0.16$ hectares, with a minimum of $7\text{m}$ meters for linear structures. Additionally, it represents LU classes, where the same LC class could be associated with different semantic classes, such as the fine-grained classes “Commercial airports”, “Sport flight airports and heliports”, and “Military airports”. We considered the ELU dataset corresponding to the Military Grid Reference System (MGRS) Sentinel-2 (S2) granule 32TPQ ($10980\times 10980$ pixels, for a total area of $110\times 110$ $km^{2}$). In the analyzed granule, the four hierarchical levels show $5$, $15$, $40$, and $90$ classes at the macro, coarse-grained, intermediate, and fine-grained levels, respectively. Please note that such an elevated number of fine-grained classes is not commonly approached in Earth Observation (EO) classification problems, where the majority of studies focus on the intermediate or coarse-grained classification task. Nonetheless, this level of precision allows a better monitoring of the class of interest and the extraction of information-rich content. Due to the high number of classes and, consequently, the large size of the hierarchical tree, we refer to this hierarchy as a complex hierarchy. A semicircular tree representing the hierarchical structure of the dataset is shown in Fig. 4.
IV-C S2 Preprocessing and Dataset Preparation
IV-C1 VHR dataset
@@ IV-C2 Multitemporal MS datasets
The input data consists of S2 multispectral (MS) satellite images provided in the Copernicus EO program. The S2 acquisitions considered have been preprocessed at Level-2A (L2A) and re-projected at $10\text{m}$/pixel resolution through a nearest neighbor approach. In the experiments, all the bands except the $60\text{m}$ resolution band were used. Due to the high-complexity of both classification problems, we collected a time series (TS) of S2 images, fully exploiting the temporal signature and allowing the detection of seasonal targets. The S2 TS considered for the Emilia Dataset consists of acquisitions ranging from January $1^{\text{st}}$ $2020$ to December $31^{\text{st}}$ $2020$. The S2 TS has been processed to achieve a more regular TS, resulting in $12$ medoid monthly composites [13].
@@
To extract a balanced training set, we considered an approach based on the heterogeneity of the input training data. In particular, we divided the initial S2 monthly composite images into non-overlapping tiles of $1000\times 1000$ pixels ($10km^{2}$). After the initial subdivision, we randomly subdivided the tiles into training, validation, and test sets by considering a split-ratio of $0.7$, $0.15$ and $0.15$, respectively. We extracted $N=2000$ representative points for each tile and class in the analyzed classification scheme. In the case of minoritarian classes having less than $N$ samples, all the available samples were extracted. The training set was generated by extracting patches of $9\times 9$ pixels around the selected points in the training tiles. To minimize spatial autocorrelation between the training patches, the corresponding reference data was extracted only for the central $3\times 3$ pixels. In this way, we collected a training dataset representative of the entirety of the classification scheme, corresponding to $\sim 1.17\%$ ($1.413.522$ pixels) of the total number of pixels available in the MGRS S2 granule.
IV-D Design of Experiments
Validation has been conducted on both datasets considering different multigranularity approaches. The reported results include the values of Overall Accuracy (OA) and mean F1 Score (mF1) for both tasks, while for the multispectral dataset also Top-3 and Top-5 metrics are reported. We considered different multigranularity modules as comparison, all built on the same backbone architecture for fairness on all the datasets. We considered the following approaches:
  1. 1.
    Baseline: the baseline with only the fine-grained classification through standard CCE minimization. Selected to evaluate the improvement gained by incorporating the hierarchy.
  2. 2.
    HAF [15]: selected as a regularization approach adopting soft labels and geometric constraints that use auxiliary classifiers at each level to ensure consistent predictions with the hierarchy, specifically aiming to reduce the severity of mistakes.
  3. 3.
    HRN [8]: included as state-of-the-art benchmark that facilitates hierarchical feature interaction by forcing finer classes to inherit attributes from coarser classes through residual skip connections.
  4. 4.
    MGM [31]: included to assess the effectiveness of using knowledge-guided relation graphs and bi-directional interaction structures that enhance coarse-grained features through gradient truncations.
  5. 5.
    HiCervix [4]: chosen as benchmark for methods that model semantic correlations within an explicit tree structure using soft labels and recursive probability summation across hierarchical paths.
  6. 6.
    ALGS [7]: selected to evaluate the impact of a global semantic distance-based penalty term for adjusting class predictions, ensuring hierarchical consistency when classification is enforced at every level. Specifically designed for RS tasks.
  7. 7.
    HCL [9]: selected as comparison for representation-based learning that disentangles class-specific embeddings at multiple levels to align them with their semnantic relationships. Specifically designed for RS tasks.
For the VHR task, we considered a ResNet50 backbone pretrained with ImageNet weights [22]. This choice is due to its deeply stacked convolutional layers and residual connections, making it well-suited for extracting highly localized, fine-grained spatial and morphological features necessary for VHR analysis. For the MS Satellite Image Time Series (SITS) classification task, we adopted a different backbone architecture designed to process multispectral and temporal information. This backbone consist in an adaptation to the RS scenario of a Swin Transformer [16] and is trained from scratch. For a detailed description of the layer topology, please refer to the cited paper [16]. The choice of this backbone is due to its shifted-window self-attention mechanism that efficiently captures both local details and long-range global contextual dependencies, making it ideal for processing complex multi-scale structural variations.
@@ IV-D Design of Experiments
We adopted the AdamW optimizer for all the architectures and configurations. The metaparameters of each different configuration have been validated considering a Tree-structured Parzen Estimator (TPE) algorithm to efficiently explore the metaparameters space. The learning rate and the weight decay were sampled from log-uniform distributions $U_{\log}([10^{-6},10^{-3}])$. We considered 50 training epochs on both datasets, together with Stochastic Weight Averaging (SWA) to enhance model generalization [19]. Considering the proposed Semantics-Aware Hierarchical Consensus (SAHC) approach, the weight $\lambda_{HC}$ associated to the HC loss has been sampled from a log-uniform distribution $U_{\log}([10^{-2},10^{2}])$, while the weights $\lambda^{h}$ have been set to $0.5$, $0.2$, and $0.3$ for the VHR dataset, and to $0.4$, $0.2$, $0.1$, and $0.3$ for the MS dataset, to prioritize the fine-grained classification and reduce the influence of the intermediate, coarse-grained, and macro levels. We implemented a dynamic consistency warmup schedule for the HC loss weight, $\lambda_{HC}$. Specifically, during the training phase, $\lambda_{HC}$ was initially fixed to 0 for the first 5 epochs to allow for stable feature extraction. Subsequently, the weight value was linearly ramped up from epoch 5 to 15 until reaching its sampled maximum value, where it remained constant for the duration of the training.
@@ V Experimental Results
TABLE I: Comparison of the Accuracies obtained by different hierarchical methodologies on the NWPU-RESISC45 dataset. The best results are reported in bold, second best are underlined
Metric Hierarchy Baseline [22] HRN [8] HAF [15] ALGS [7] HCL [9] MGM [31] HiCervix [4] SAHC
OA (%) Fine-grained 96.08 96.48 96.16 96.32 95.87 96.11 96.10 96.63
Intermediate 97.38 97.30 97.35 97.16 97.24 97.19 97.56
Coarse-grained 97.86 97.86 97.78 97.73 97.84 97.70 97.98
mF1 (%) Fine-grained 96.08 96.47 96.15 96.33 95.85 96.10 96.09 96.63
Intermediate 96.74 96.37 96.65 96.26 96.46 96.37 96.89
Coarse-grained 97.56 97.52 97.47 97.36 97.41 97.37 97.67
Params (M) 25.8 35.5 25.9 25.9 27.0 35.5 25.9 25.9
FLOP (T) 6.22 7.03 6.22 6.22 6.51 6.88 6.22 6.22
The experiments focus on validating the proposed SAHC approach over different datasets, comparing it to other state-of-the-art hierarchical approaches on different levels of the hierarchy. We first focus on the results on the specific tasks, i.e., scene classification on the VHR dataset and semantic segmentation of MS SITS. Then, we analyze the importance of the hierarchical weights selection $\lambda^{h}$ on the NWPU dataset. Finally, we discuss a visualization of the trained hierarchy projection matrices to identify the key hierarchical relationships learnt by the proposed approach.